OpenStudy (anonymous):

PLEASE HELP ME AND WILL GIVE MEDAL! 2. Justify each step of the proof. Conjecture: If 2(x2 + 5) = 60, then x is 5 or –5. 1. 2(x2 + 5) = 60 1. 2. 2x2 + 10 = 60 2. 3. 2x2 = 50 3. 4. x2 = 25 4. 5. x = ±5 5.

OpenStudy (anonymous):

@cram

OpenStudy (anonymous):

@Nurali pls help anyone

OpenStudy (nurali):

>> There's no justification needed for step 1; >> Step 2 is true because of the distribuitive property of multiplication: >> 2x^2 + 10 is the same as 2(x^2 + 5) >> Step 3 is true because of the commutative property of addition, and the >> existence of a right identity for addition: effectively, you've added -10 to >> both sides, as in 2^2 + 10 - 10 = 60 - 10 giving 2^2 = 50 >> Step 4 is justified by the commutative property of multiplication; you're >> multiplying each side of the equation by 1/2 >> Step 5 is justified by the argument that no other number quite fits.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

@Nurali did you get my pm?

OpenStudy (nurali):

Anytime.